# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v5

@@ Abstract
As a common image editing operation, image composition aims to combine the foreground from one image and another background image, resulting in a composite image. However, there are many issues that could make the composite images unrealistic. These issues can be summarized as the inconsistency between foreground and background, which includes appearance inconsistency (e.g., incompatible illumination), geometry inconsistency (e.g., unreasonable size), and semantic inconsistency (e.g., mismatched semantic context). Image composition task could be decomposed into multiple sub-tasks, in which each sub-task targets at one or more issues. Specifically, object placement aims to find reasonable scale, location, and shape for the foreground. Image blending aims to address the unnatural boundary between foreground and background. Image harmonization aims to adjust the illumination statistics of foreground. Shadow generation aims to generate plausible shadow for the foreground. These sub-tasks can be executed sequentially or parallelly to acquire realistic composite images. To the best of our knowledge, there is no previous survey on image composition. In this paper, we conduct comprehensive survey over the sub-tasks and combinatorial task of image composition. For each one, we summarize the existing methods, available datasets, and common evaluation metrics. Datasets and codes for image composition are summarized at https://github.com/bcmi/Awesome-Image-Composition. We have also contributed the first image composition toolbox: libcom https://github.com/bcmi/libcom, which assembles 10+ image composition related functions (e.g., image blending, image harmonization, object placement, shadow generation, generative composition). The ultimate goal of this toolbox is solving all the problems related to image composition with simple ‘import libcom’.
@@ I Introduction
+ + $\cdot$ $\cdot$ + [183, 141, 205, 16, 191, 187, 142]
The appearance inconsistency is including but not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [78]. For the first issue, the foreground is usually extracted using image segmentation [108] or matting [180, 33] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [172, 198] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could be inharmonious (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [156, 22, 18] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the whole image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [92, 203, 134] or reflection generation [105] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. As far as we are concerned, there are only few works [105, 171] on generating reflection for the inserted object probably due to the limited application scenarios, so we focus on shadow generation in this paper. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [168], deblurring [196], denoising [153] techniques.
@@
So far, we have introduced several sub-tasks (e.g., image blending, image harmonization, shadow generation, object placement), in which one sub-task targets at one or multiple issues. Previous works usually focus on one sub-task or perform multiple sub-tasks sequentially (i.e., image blending followed by image harmonization) as shown in Fig. 3(a). The reasonable sequential order is as follows. Given a pair of foreground and background, we first use object placement to find suitable scale and location for the foreground, and use image blending to refine the boundary between foreground and background. Then, we use image harmonization to adjust the foreground illumination and shadow generation to generate plausible shadow for the foreground. Recently, as the diffusion models have demonstrated unprecedented generation ability, some works [183, 141] utilize diffusion models to perform multiple sub-tasks (e.g., image blending, image harmonization, view synthesis) parallelly as shown in Fig. 3(b). Given a pair of foreground and background with bounding box, they propose one unified model to directly produce the composite image, in which the foreground is blended seamlessly and harmoniously into the background. These methods re-generate the foreground object instead of making restrained adjustments for the foreground object, so we refer to them as generative image composition methods. We summarize all the potential issues and the corresponding methods to solve them in Table I. In the last row, we use $\cdot$ instead of + for “shadow” and “geometry”, because the shadow and geometry issues are only partially solved, which will be explained in Section VI. Another problem for parallel pipeline is that whether we can switch on or switch off some subtasks when performing them parallelly, with the goal of more controllable image composition. Based on this idea, Zhang et al. [191] built a prototype named ControlCom, which uses a 2-dim binary indicator vector to selectively adjust the illumination or pose of foreground object. In the ideal case, assuming that we have in total $K$ subtasks, a $K$-dim binary indicator vector could be used to indicate the selected subtasks.
@@
  • To the best of our knowledge, this is the first comprehensive survey on deep image composition.
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [154, 197, 88, 81, 80] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [154] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts the affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [188] adopted spatial transformer network (STN) [58] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [197] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensure the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [208] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [202] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Azadi et al. [2] employed STN to warp the foreground and relative appearance flow network to change the viewpoint of foreground. Additionally, they investigated on self-consistency constraint, that is, the generated composite image could be decomposed back to the foreground and background. ST-GAN [88] proposed to warp a foreground object into a background image with iterative spatial transformations predicted by STN. As a follow-up work, Kikuchi et al. [65] replaced the iterative spatial transformations in [88] with one-shot spatial transformation.
@@ II-C Datasets and Evaluation Metrics
In some previous works [154, 29], object placement is used as data augmentation strategy to facilitate the downstream tasks (e.g., object detection, instance segmentation). Therefore, they make use of existing object detection and instance segmentation datasets [89, 28, 21, 38]. In particular, the foregrounds are cropped out based on the annotated segmentation masks. After removing the foreground objects, the remaining incomplete background images are restored to complete background images by using image inpainting techniques [185, 93, 186]. In this manner, triplets of foregrounds, backgrounds, and ground-truth composite images can be obtained. Some other works focus on specific applications like 2D virtual try-on [88, 65, 81] (e.g., placing glasses/hats on human faces) or logo composition [80] (e.g., attaching logo to product image), so they need to collect foregrounds and backgrounds specifically for these applications. More recently, Liu et al. [94] released a large-scale object placement assessment dataset named OPA, which consists of 73,470 composite images and their binary rationality labels. OPA dataset is constructed by compositing the foregrounds and backgrounds from COCO dataset [89], followed by manually labelling the rationality of obtained composite images. A large number of annotated composite images could greatly facilitate the research on object placement.
@@ IV-A Rendering based Methods
Conventional image relighting [117, 137, 167, 199, 152] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [117, 199], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, and 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption for the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [156, 18, 20, 97, 102, 146, 14] using paired supervision have been developed. Tsai et al. [156] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [139] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [22] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [49] employed self-attention [165] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [18] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [19] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [19] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those apparently inharmonious composite images. Inspired by [19], Valanarasu et al. [157] proposed to extract style code from part of background.
In [90], they reframed image harmonization as a background-to-foreground style transfer problem and introduced region-aware adaptive instance normalization (AdaIn) to transfer visual style from the background to the foreground. A succeeding work [48] extended [90] by searching foreground-relevant background regions and transferring foreground-relevant style from background to foreground. They also extended the triplet loss in [19] to contrastive loss.
@@
Inspired by traditional image harmonization methods [182, 143] which applied color transformation to adjust the foreground appearance, Cong et al. [20] proposed to learn color transformation using deep learning for image harmonization. They combined color-to-color transformation and pixel-to-pixel transformation in a unified framework coherently. Several other works [181, 86, 127, 64, 43, 161] also proposed to predict various types of color transformations (e.g., color filter, rendering curve, linear transformation) for efficient image harmonization. Beyond different color transformations, some works [179, 146] explore different color spaces.
Some recent works [113, 133, 13] concurred that dynamic kernels acting upon feature maps can boost the harmonization performance. Furthermore, [113, 133] pointed out the importance of global information in dynamic kernel prediction. Niu et al. [112] studied domain adaptive image harmonization by treating different datasets as different domains. Specifically, an automatic augmentation network was developed to enrich the illumination diversity of a target domain with limited data.
@@ IV-C Variants of Image Harmonization Task
The existing painterly image harmonization methods [104, 119, 10, 99, 166, 115, 114] can be roughly categorized into optimization-based methods and feed-forward methods. Optimization-based methods optimize the input image to minimize the style loss and content loss, which is very time-consuming. For example, Luan et al. [104] proposed to optimize the input image with two passes, in which the first pass aims at robust coarse harmonization and the second pass targets at high-quality refinement. Feed-forward methods send the input image through the model to output the harmonized result. For example, Peng et al. [119] applied adaptive instance normalization to match the means and variances between the feature map of composite image and that of artistic background. Cao et al. [10] performed painterly image harmonization in both frequency domain and spatial domain, considering that artistic paintings often have periodic textures and patterns which appear regularly. Lu et al. [99] introduced diffusion model to painterly image harmonization, which can significantly outperform GAN-based methods when the background has dense textures or abstract style. Niu et al. [115] divided styles into low-level styles (e.g., color, simple pattern) and high-level styles (e.g., complex pattern), and devised a progressive network which can harmonize a composite image from low-level styles to high-level styles progressively. Niu et al. [114] proposed style-level supervision based on pairs of artistic objects and photographic objects, considering that it is hard to obtain pixel-wise supervision based on pairs of artistic images and photographic images. Niu et al. [114] also contributed an artistic object dataset which contains the segmentation masks and similar photographic objects for artistic objects.
@@ IV-E Datasets and Evaluation Metrics
Replacement: A natural way to build image harmonization dataset is collecting a set of foreground images captured in different illumination conditions, followed by replacing one foreground with another counterpart. For example, Transient Attributes Database [70] contains 101 sets, in which each set has well-aligned images for the same scene captured in different conditions (e.g., weather, time of the day, season). This dataset has been used to construct pairs in Hday2night sub-dataset in iHarmony4 [18]. However, collecting the dataset like [70] calls for capturing the same scene with a fixed camera for a long time, which is hard to be realized in practice. To obtain the foregrounds in different capture conditions, Song et al. [140] proposed an interesting way to construct GMS Dataset. Specifically, they place the same physical model (3D foreground object) in different lighting conditions to capture different images and align the foregrounds in different images. Nevertheless, the collection cost is still very high and the diversity of foreground is restricted. In lieu of capturing images in the real world, an easy alternative is varying the illumination condition in a virtual environment. Cao et al. [9] constructed RdHarmony dataset by varying the lighting condition of the same scene using 3D rendering techniques. Within a set of images with the same scene yet various lighting conditions, the composite images could be obtained by exchanging the foregrounds between two images. Similarly, Guo et al. [45] constructed HVIDIT dataset based on the rendered dataset [27]. However, the rendered images have large domain gap with real images, so the harmonization model trained on rendered images cannot be directly applied to real test images.

### v6

@@ Abstract
As a common image editing operation, image composition (object insertion) aims to combine the foreground from one image and another background image, resulting in a composite image. However, there are many issues that could make the composite images unrealistic. These issues can be summarized as the inconsistency between foreground and background, which includes appearance inconsistency (e.g., incompatible illumination), geometry inconsistency (e.g., unreasonable size), and semantic inconsistency (e.g., mismatched semantic context). Image composition task could be decomposed into multiple sub-tasks, in which each sub-task targets at one or more issues. Specifically, object placement aims to find reasonable scale, location, and shape for the foreground. Image blending aims to address the unnatural boundary between foreground and background. Image harmonization aims to adjust the illumination statistics of foreground. Shadow (resp., reflection) generation aims to generate plausible shadow (resp., reflection) for the foreground. These sub-tasks can be executed sequentially or parallelly to acquire realistic composite images. To the best of our knowledge, there is no previous survey on image composition (object insertion). In this paper, we conduct comprehensive survey over the sub-tasks and combinatorial task of image composition (object insertion). For each one, we summarize the existing methods, available datasets, and common evaluation metrics. Datasets and codes for image composition are summarized at https://github.com/bcmi/Awesome-Image-Composition. We have also contributed the first image composition toolbox: libcom https://github.com/bcmi/libcom, which assembles 10+ image composition related functions (e.g., image blending, image harmonization, object placement, shadow generation, generative composition). The ultimate goal of this toolbox is solving all the problems related to image composition with simple ‘import libcom’.
@@ I Introduction
+ + + [80]
+ + $\cdot$ $\cdot$ + [202, 152, 227, 213, 209, 153]
+ + + + [53]
+ + + $\cdot$ + [190]
+ + + + + [230]
+ + + + + + [86]
The appearance inconsistency is including but not limited to: 1) unnatural boundary between foreground and background; 2) incompatible illumination statistics between foreground and background; 3) missing or implausible shadow and reflection of foreground; 4) resolution, sharpness, and noise discrepancy between foreground and background [85]. For the first issue, the foreground is usually extracted using image segmentation [118] or matting [199, 36] algorithms. However, the foregrounds may not be precisely delineated, especially at the boundaries. When pasting the foreground with jagged boundaries on the background, there would be obvious color artifacts along the boundary. To solve this issue, image blending [191, 220] aims to address the unnatural boundary between foreground and background, so that the foreground could be seamlessly blended with the background. For the second issue, since the foreground and background may be captured in different conditions (e.g., weather, season, time of the day, camera setting), the obtained composite image could be inharmonious (e.g., foreground captured in the daytime and background captured at night). To solve this issue, image harmonization [170, 25, 21] aims to adjust the illumination statistics of foreground to make it more compatible with the background, so that the whole image looks more harmonious. For the third issue, when pasting the foreground on the background, the foreground may also affect the background with shadow or reflection. To solve this issue, shadow generation [101, 225, 145] or reflection generation [114] focus on generating plausible shadow or reflection for the foreground according to both foreground and background information. As far as we are concerned, there are only few works [114, 189, 165] on generating reflection for the inserted object probably due to the limited application scenarios, so we focus on shadow generation in this paper. For the fourth issue, the foreground and background may be from two images with different resolutions, blur degrees, and noise patterns. The resolution (resp., sharpness, noise) discrepancy between them could be mitigated by using super-resolution [186], deblurring [218], denoising [167] techniques.
@@
So far, we have introduced several sub-tasks (e.g., image blending, image harmonization, shadow generation, object placement), in which one sub-task targets at one or multiple issues. Previous works usually focus on one sub-task or perform multiple sub-tasks sequentially (i.e., image blending followed by image harmonization) as shown in Fig. 3(a). The reasonable sequential order is as follows. Given a pair of foreground and background, we first use object placement to find suitable scale and location for the foreground, and use image blending to refine the boundary between foreground and background. Then, we use image harmonization to adjust the foreground illumination and shadow generation to generate plausible shadow for the foreground. Recently, as the diffusion models have demonstrated unprecedented generation ability, some works [202, 152] utilize diffusion models to perform multiple sub-tasks (e.g., image blending, image harmonization, view synthesis) parallelly as shown in Fig. 3(b). Given a pair of foreground and background with bounding box, they propose one unified model to directly produce the composite image, in which the foreground is blended seamlessly and harmoniously into the background. These methods re-generate the foreground object instead of making restrained adjustments for the foreground object, so we refer to them as generative image composition methods. We summarize all the potential issues and the corresponding methods to solve them in Table I. Note that we use $\cdot$ instead of + in some slots, because these issues are only partially solved. Another problem for parallel pipeline is that whether we can switch on or switch off some subtasks when performing them parallelly, with the goal of more controllable image composition. Based on this idea, Zhang et al. [213] built a prototype named ControlCom, which uses a 2-dim binary indicator vector to selectively adjust the illumination or pose of foreground object. In the ideal case, assuming that we have in total $K$ subtasks, a $K$-dim binary indicator vector could be used to indicate the selected subtasks.
@@
  • To the best of our knowledge, this is the first comprehensive survey on deep image composition (object insertion).
@@ II-B2 Instance-specific Object Placement
Generative approaches: Generative approaches [168, 219, 97, 90, 89, 204, 182] predict different types of spatial transformations (e.g., shifting and scaling, affine transformation, perspective transformation) for the foreground object, which is more flexible and powerful than category-specific object placement methods. For instance, Tripathi et al. [168] developed a model with generator, discriminator, and target network. Given a pair of background and foreground, the generator predicts the affine transformation for the foreground object to produce a composite image. The produced composite image is expected to fool the discriminator and fit the target network corresponding to a downstream task (e.g., object detection). Similarly, Zhan et al. [210] adopted spatial transformer network (STN) [63] to predict the warping parameters under an adversarial learning framework. To produce multiple reasonable placements, Zhang et al. [219] combined the foreground feature, background feature, and a random vector to predict the object placement. Moreover, they ensure the diversity of object placement by enforcing the pairwise distances between predicted placements to approach those between corresponding random vectors. To promote the diversity of generated placements, Zhou et al. [234] established the bijection between random vector and positive composite image. Moreover, they reformulated object placement as a graph completion task. In particular, background nodes have both content features and placements, while the inserted foreground node only has content feature, giving rise to an incomplete graph. Hence, they estimated the missing placement for the foreground node to complete the graph. Zhang et al. [224] proposed to make sequential decisions to produce a reasonable placement by using reinforcement learning. Qin et al. [134] employed a pre-trained large multi-modal model to generate a caption containing the placement information, and then predicted the placement bounding boxes.
Azadi et al. [2] employed STN to warp the foreground and relative appearance flow network to change the viewpoint of foreground. Additionally, they investigated on self-consistency constraint, that is, the generated composite image could be decomposed back to the foreground and background. ST-GAN [97] proposed to warp a foreground object into a background image with iterative spatial transformations predicted by STN. As a follow-up work, Kikuchi et al. [71] replaced the iterative spatial transformations in [97] with one-shot spatial transformation.
@@ II-C Datasets and Evaluation Metrics
In some previous works [168, 32], object placement is used as data augmentation strategy to facilitate the downstream tasks (e.g., object detection, instance segmentation). Therefore, they make use of existing object detection and instance segmentation datasets [98, 31, 24, 41]. In particular, the foregrounds are cropped out based on the annotated segmentation masks. After removing the foreground objects, the remaining incomplete background images are restored to complete background images by using image inpainting techniques [205, 102, 207]. In this manner, triplets of foregrounds, backgrounds, and ground-truth composite images can be obtained. Some other works focus on specific applications like 2D virtual try-on [97, 71, 90] (e.g., placing glasses/hats on human faces) or logo composition [89] (e.g., attaching logo to product image), so they need to collect foregrounds and backgrounds specifically for these applications. More recently, Liu et al. [103] released a large-scale object placement assessment dataset named OPA, which consists of 73,470 composite images and their binary rationality labels. OPA dataset is constructed by compositing the foregrounds and backgrounds from COCO dataset [98], followed by manually labelling the rationality of obtained composite images. A large number of annotated composite images could greatly facilitate the research on object placement. Qin et al. [134] established the OPAZ dataset following the format of OPA.
@@ IV-A Rendering based Methods
Conventional image relighting [127, 148, 185, 221, 166] aims to adjust the appearance of an image or the object in an image as lit by novel illumination. With some adaptation, image relighting can also be used to adjust the foreground appearance according to the illumination of a new background [127, 221, 11, 59], which bears some resemblance to image harmonization. However, they usually achieve this goal by inferring explicit illumination condition, material properties, and 3D geometry, in which the supervision for these information is difficult and expensive to acquire. Besides, they generally have strong assumption for the light source, which may not generalize well to complicated real-world scenes.
@@ IV-B Non-rendering based Methods
With the emergence of image harmonization datasets consisting of paired training data (see Section IV-E), abundant image harmonization methods [170, 21, 23, 106, 111, 158, 16, 206, 184, 130] using paired supervision have been developed. Tsai et al. [170] proposed the first end-to-end CNN network for image harmonization and leveraged auxiliary semantic segmentation branch to enhance the basic image harmonization network. Another work Sofiiuk et al. [150] also utilized high-level semantic features, which are inserted into the encoder to provide auxiliary information. Cun and Pun [25] designed an additional Spatial-Separated Attention Module to deal with foreground and background feature maps separately. Hao et al. [52] employed self-attention [180] mechanism to propagate relevant information from background to foreground. By treating different capture conditions as different domains, Cong et al. [21] proposed a domain verification discriminator to pull close the foreground domain and background domain. Similarly, Cong et al. [22] formulated image harmonization as background-guided domain translation task, in which the domain code of background is directly used to guide the harmonization process. One byproduct of [22] is predicting the inharmony level of an image by comparing the domain codes of foreground and background, so that we can selectively harmonize those apparently inharmonious composite images. Inspired by [22], Valanarasu et al. [171] proposed to extract style code from part of background.
In [99], they reframed image harmonization as a background-to-foreground style transfer problem and introduced region-aware adaptive instance normalization (AdaIn) to transfer visual style from the background to the foreground. A succeeding work [51] extended [99] by searching foreground-relevant background regions and transferring foreground-relevant style from background to foreground. They also extended the triplet loss in [22] to contrastive loss. Some subsequent works [66, 15] adopted the similar idea of searching the background regions matching the foreground region.
@@
Inspired by traditional image harmonization methods [201, 155] which applied color transformation to adjust the foreground appearance, Cong et al. [23] proposed to learn color transformation using deep learning for image harmonization. They combined color-to-color transformation and pixel-to-pixel transformation in a unified framework coherently. Several other works [200, 95, 139, 70, 46, 175, 117] also proposed to predict various types of color transformations (e.g., color filter, rendering curve, linear transformation) for efficient image harmonization. Beyond different color transformations, some works [198, 158] explored different color spaces.
Some works [123, 144, 14] concurred that dynamic kernels acting upon feature maps can boost the harmonization performance. Furthermore, [123, 144] pointed out the importance of global information in dynamic kernel prediction. Niu et al. [122] studied domain adaptive image harmonization by treating different datasets as different domains. Specifically, an automatic augmentation network was developed to enrich the illumination diversity of a target domain with limited data.
Recently, some diffusion-based image harmonization models [92, 17, 233, 232, 138] have applied conditional diffusion model to image harmonization task. Zhou et al. [232] proposed to modify VAE decoder to alleviate the image distortion issue of diffusion model. Ren et al. [138] proposed to inject background illumination information into diffusion model.
@@ IV-C Variants of Image Harmonization Task
The existing painterly image harmonization methods [113, 129, 10, 108, 181, 125, 124] can be roughly categorized into optimization-based methods and feed-forward methods. Optimization-based methods optimize the input image to minimize the style loss and content loss, which is very time-consuming. For example, Luan et al. [113] proposed to optimize the input image with two passes, in which the first pass aims at robust coarse harmonization and the second pass targets at high-quality refinement. Li et al. [87] proposed to optimize the latent features of diffusion model based on content loss and style loss.
Feed-forward methods send the input image through the model to output the harmonized result. For example, Peng et al. [129] applied adaptive instance normalization to match the means and variances between the feature map of composite image and that of artistic background. Cao et al. [10] performed painterly image harmonization in both frequency domain and spatial domain, considering that artistic paintings often have periodic textures and patterns which appear regularly. Niu et al. [125] divided styles into low-level styles (e.g., color, simple pattern) and high-level styles (e.g., complex pattern), and devised a progressive network which can harmonize a composite image from low-level styles to high-level styles progressively. Niu et al. [124] proposed style-level supervision based on pairs of artistic objects and photographic objects, considering that it is hard to obtain pixel-wise supervision based on pairs of artistic images and photographic images. Niu et al. [124] also contributed an artistic object dataset which contains the segmentation masks and similar photographic objects for artistic objects. Sun and Zhang [154] applied dynamic kernel to painterly harmonization. Lu et al. [108] is the first work introducing diffusion model to painterly image harmonization, which can significantly outperform GAN-based methods when the background has dense textures or abstract style.
@@ IV-E Datasets and Evaluation Metrics
Replacement: A natural way to build image harmonization dataset is collecting a set of foreground images captured in different illumination conditions, followed by replacing one foreground with another counterpart. For example, Transient Attributes Database [76] contains 101 sets, in which each set has well-aligned images for the same scene captured in different conditions (e.g., weather, time of the day, season). This dataset has been used to construct pairs in Hday2night sub-dataset in iHarmony4 [21]. However, collecting the dataset like [76] calls for capturing the same scene with a fixed camera for a long time, which is hard to be realized in practice. To obtain the foregrounds in different capture conditions, Song et al. [151] proposed an interesting way to construct GMS Dataset. Specifically, they place the same physical model (3D foreground object) in different lighting conditions to capture different images and align the foregrounds in different images. Nevertheless, the collection cost is still very high and the diversity of foreground is restricted. In lieu of capturing images in the real world, an easy alternative is varying the illumination condition in a virtual environment. Cao et al. [9] constructed RdHarmony dataset by varying the lighting condition of the same scene using 3D rendering techniques. Within a set of images with the same scene yet various lighting conditions, the composite images could be obtained by exchanging the foregrounds between two images. Similarly, Guo et al. [48] constructed HVIDIT dataset based on the rendered dataset [30]. [60, 59] rendered the same 3D foreground model using different illumination maps. However, the rendered images have large domain gap with real images, so the harmonization model trained on rendered images cannot be directly applied to real test images.